write the equation of a quadratic function with roots -1 and 1 and a vertex at (0,-3)
Answer by nerdybill(7384)

write the equation of a quadratic function with roots -1 and 1 and a vertex at (0,-3)
.
Vertex form of a quadratic:
y = a(x-h)^2 + k
Since the problem already gives you the vertex at (0,-3) we now have:
y = a(x-0)^2 - 3
y = ax^2 - 3
To find 'a', we use one of the roots (1,0) and substitute it in for (x,y):
0 = a(1)^2-3
0 = a-3
3 = a
.
Now we have:
y = 3x^2 - 3 (answer in vertex form)